OpenStudy (anonymous):

Which is the solution set for –7 ≤ 4x + 1 ≤ 5?

OpenStudy (anonymous):

x ≥ and x ≤ B. x ≥ or x ≤ C. x ≤ –2 or x ≥ 1 D. x ≥ –2 and x ≤ 1

OpenStudy (anonymous):

–7 ≤ 4x + 1 ≤ 5 First subtract 1 from each portion of the inequality.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

–8 ≤ 4x ≤ 4 next divide each portion by 4

OpenStudy (anonymous):

ah i know i get it. thanks

OpenStudy (anonymous):

Good. Just make sure you also know that if you have to multiply or divide by a NEGATIVE during these simplification then you have to reverse the directions of the inequalities. \[-6\le -2x \le 8\rightarrow 3\ge x \ge -4\]

OpenStudy (anonymous):

oh ok. thanks!